OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6. is it a. y = -1/3x + 4 b. y = 3x + 20 c. y = -3x - 16 d. none of the above

OpenStudy (anonymous):

.perpendicular to y = 3x + 6. perpendicular to a line means, its slope m = -1/m of the line perpendicular to it m=-1/3 and at point (6, 2) use y-y1=m(x-x1) y-2=(-1/3)(x-6) y-2=-1/3 x + 2 y=-1/3 x +4